The stockholders’ equity section of Fauberg Marigny Corporation at December 31 is as follows.

FAUBERG MARIGNY CORPORATION
Balance Sheet (partial)
Stockholders' equity
Paid-in capital
Preferred stock, cumulative, 10,000 shares authorized,
    5,000 shares issued and outstanding
$300,000
Common stock, no par, 750,000 shares authorized,
    300,000 shares issued
1,500,000
Total paid-in capital 1,800,000
Retained earnings 2,050,000
Total paid-in capital and retained earnings 3,850,000
Less: Treasury stock (5,000 common shares) (64,000 )
Total stockholders' equity $3,786,000


From a review of the stockholders’ equity section, answer the following questions.

1)How many shares of common stock are outstanding?

2) Assuming there is a stated value, what is the stated value of the common stock?Stated value of common stock per share.

3)What is the par value of the preferred stock?

Par value of preferred stock

per share

4)

If the annual dividend on preferred stock is $18,000, what is the dividend rate on preferred stock?

Dividend rate on preferred stock

5)

If dividends of $36,000 were in arrears on preferred stock, what would be the balance in retained earnings?

Balance for Retained Earnings
Answer #1

1) Outstanding shares = 300000-5000 = 295000 Shares

2) Stated value = 1500000/300000 = 5 per share

3) Par value of preferred stock = 300000/5000 = 60 per share

4) Dividend rate = 18000/300000 = 6%

5) Balance in retained earnings = 2050000-36000 = 2014000